# DeltaDorsal: Enhancing Hand Pose Estimation with Dorsal Features in Egocentric Views

### v1

@@ 1. Introduction
Hand tracking has become a cornerstone of modern human–computer interaction (Chakravarthi et al., 2023; Ungureanu et al., 2020; Zhang et al., 2020), enabling intuitive gesture-based control in Extended Reality (XR), wearables (Yeo et al., 2019), and other interactive systems (Pei et al., 2022; Mollyn and Harrison, 2024). As these technologies expand, the need for accurate and robust hand pose estimation continues to grow, especially from an egocentric point-of-view. However, a key challenge for egocentric hand pose estimation remains: self-occulusion. From this viewpoint, the palm often partially or fully obscures the very fingers being tracked. Our analysis of prominent egocentric hand datasets (Ohkawa et al., 2023; Grauman et al., 2024b; Fan et al., 2023; Kwon et al., 2021) reveals that a finger is fully occluded in nearly a fifth of all scenarios, with partial occlusions occurring even more frequently.
@@
    An evaluation of DeltaDorsal’s performance on 12 participants versus state-of-the-art baselines, as well as analyses with respect to occlusion, skin tone, image size, and backbone.
@@
    Several exemplary applications of DeltaDorsal in key hand interactions, including pinching, tapping, and isometric force click.
@@ 2.2. Hand Pose Estimation Leveraging Skin Features
Vision-based approaches observing skin features free the dorsum area from on-hand sensors. Opisthenar uses a wrist-worn camera to capture dorsal features for gesture and tap recognition (Yeo et al., 2019). DorsalNet uses the same camera platform with a hand simulator to capture motion images of the dorsum to predict 3D hand poses. These camera-based methods all require hardware setups on the wrist or arm, making them less usable compared to the more traditional egocentric hand tracking systems.
@@ 3.1. Measuring Self-Occlusion
Using the hand mesh $\Theta$, we project vertices $V$ into the egocentric camera space using the camera extrinsic parameters $V_{\textrm{cam}}=RV_{\textrm{world}}+t$. For each face of the MANO model, we compute its normal vector and filter to only faces with a normal vector facing the camera (z-z). We then apply Z-buffering (Catmull, 1978) using this filtered list of faces. The 3D mesh is projected into the camera’s image plane where each mesh face becomes a 2D triangle. We then create an array of depth values, known as a Z-buffer, to track the closest distance to some face at that pixel. For each pixel, we compute the triangle membership through barycentric coordinate sampling and track the interpolated depth at that point. If the interpolated depth is equal or less than the stored Z-buffer depth, the face is marked as visible; otherwise it is occluded and ignored. We then compute the surface area of each visible face as a measurement of mesh visibility. This measurement only considers visibility in respect to other faces of the hand and the camera view. Thus, this only measures the amount of self-occlusion per frame and does not account for occlusion from manipulating other objects or body parts, which disproportionately affects the fingers and palm.
@@ 3.2. Occlusion Prevalence in Egocentric Data
Hand self-occlusion. As shown in Figure 3a, more than 20% of frames across all evaluated datasets exhibit at least one occluded finger, with over 5% containing two or more occluded fingers. These values are a conservative estimate as we do not consider any articulating objects, which are far more likely to occlude the fingers than the dorsum of the hand. These self-occlusion cases often occurs because of three primary scenarios: The hand is angled downwards such that the back of the hand occludes the fingers, the hand is angled out such that the index finger occludes the other fingers, or the hand is partially out of frame. Our analysis indicates that there is a sizable proportion of egocentric hand capture data that is affected by self-occlusion which follows findings from Fan et al. (Fan et al., 2024).
Hand dorsum visibility. We then analyze how much of the dorsum is visible in these situations by computing the distribution of dorsal face visibility with at least one finger over 90% occluded. We find that for every dataset, the median dorsal visibility is above 40% and up to 80%, indicating that that a sizable proportion of frames with occluded fingers still have visible dorsal features. Thus, dorsal features offer an alternative rich and stable signal when hand and finger silhouettes are not available.
@@ 3.3. Impact of Occlusion on Egocentric Hand Pose Estimators
We find that for both HaMeR and HandOccNet, there is a statistically significant ($p<.0001$) negative correlation between finger visibility and error ($m=-15.84\degree$, $R^{2}=0.212$ for HaMeR, $m=-9.11\degree$, $R^{2}=0.095$ for HandOccNet). This is especially apparent in HaMeR, which has been widely regarded as a standard in pose estimation since its release and has become the backbone of many dataset annotation and hand pose estimation systems (Zhao et al., 2024a; Prakash et al., 2024). Notably for HandOccNet, we found that the model would often default to outputting a neutral hand position in self-occlusion situations, leading to serious prediction errors most likely due to the model’s focus on predicting hands occluded behind other objects. Our findings are consistent with our intuition that most existing hand tracking models rely on the general silhouette to predict pose. Often, their architecture prevents them from using skin features due to the limited token and image size allowed.
@@ 3.4. Our Proposal: Investigate Dorsal Features
Given our findings, we choose to investigate dorsal features as a key signal to improve hand pose estimation in egocentric POVs. We lay out our work as follows: we first perform a high-resolution data collection of dorsal features from different hand gestures to capture and analyze dorsal features in detail. Using this dataset, we propose and train a new model that normalizes visual features from the dorsum to a neutral hand pose, similar to how DorsalNet uses motion images (Wu et al., 2020), to isolate dorsal features even through global motion. We then evaluate our system against baseline hand pose estimation models and across skin tones and image resolutions. Finally, we demonstrate the utility of dorsal features beyond pose estimation by evaluating two key interactions, pinch and tap, and an isometric click classifier.
@@ 5.2. Network Architecture
DeltaDorsal adopts a two-stream architecture built on a transformer backbone with a lightweight convolutional change encoder followed by a regression head. A visualization of this architecture is shown in Figure 6. Using a pretrained DINOv3 Vision Transformer (ViT) (Siméoni et al., 2025) following the “large” design (ViT-L) with the last three blocks unfrozen, we first compute the dense image features for both the reference $I_{0}$ and target $I_{t}$ using the image of the dorsum of the hand, $F_{0}$ and $F_{t}$ respectively with dimensions equal to the original image size divided by the patch size ($16\times 16$). Notably, model weights are shared across each stream to ensure that $F_{0}$ and $F_{t}$ maintain the same feature space. A change encoder then processes this feature map by fusing the feature delta, $F_{t}-F_{0}$, cosine similarity per patch, $\cos(F_{t},F_{0})(x,y)$, $F_{t}$, and $F_{0}$ and feeding this data through a small convolutional network to generate the encoded features $X$. Cosine similarity features are visualized in Figure 5. By fusing both the original features, feature deltas, and cosine similarities, we allow the model to select between general image features and changes in dorsal features as needed. We settled on a dense featurizer backbone after preliminary experiments with optical flow methods like RAFT (Teed and Deng, 2020) and Gunnar-Farneback optical flow (Farnebäck, 2003) had difficulty separating skin deformations from global movements of the hand. Furthermore, traditional feature extractors like ResNet (He et al., 2015) and ConvNeXt (Liu et al., 2022) failed to generalize to out-of-distribution subjects. We believe this limitation arises because these models are not well-suited to capture the subtle soft-body surface deformations in the back of the hand which often do not demonstrate distinct, localized feature changes. Finally, a regression head applies a small convolutional neck for local refinement, pools spatial features, and passes pooled vectors through a shallow multi-layer perceptron to predict the pose vector $\beta$ ($(15,3)$ for 15 axis-angle rotations).
@@ 5.3. Implementation Details
All systems were implemented using PyTorch (Paszke et al., 2019). We used the AdamW optimizer (Loshchilov and Hutter, 2019) with a batch size of 96 for our training. During training, we randomly selected a reference frame from the corresponding calibration data capture. Please refer to our codebase and configuration files for specific parameter settings for experiments. All training and evaluations were performed using PyTorch Distributed Data Parallel on a system with an AMD Ryzen Threadripper PRO 3955WX 3.90GHz, four Nvidia RTX A5500s, and 256GB of RAM.
@@ 6.1. Pose Estimation Performance Study
Performance across occlusion levels. We also evaluate the performance of our baselines and system under different occlusion levels after fine-tuning. As shown in Figure 8, we find that our model has a smaller dropoff in performance compared to both HaMeR and HandOccNet ($m=-4.59\degree$, $R^{2}=0.048$ for ours, $m=-8.42\degree$, $R^{2}=0.117$ for HaMeR, $m=-9.11\degree$, $R^{2}=0.095$ for HandOccNet) [^1] We denote slope as $m$. HandOccNet’s general performance was not significantly changed after finetuning due to previously mentioned training difficulties. When comparing performance at 50% finger visibility and lower, DeltaDorsal outperforms HaMeR across both MPJAE (Ours: $7.59\pm 4.38\degree$, HaMeR: $9.24\pm 5.54\degree$, $17.86\%$) and PA-MPJPE (Ours: $8.47\pm 3.84$mm, HaMeR: $9.03\pm 3.65$$m$, $6.20\%$). These findings are consistent with our intuition that existing hand pose estimation models rely heavily on finger silhouettes for pose prediction. Examples of this poor performance are shown in Figure 1 and Figure 7, where HaMeR and HandOccNet both fail to accurately predict hand pose when the fingers are occluded.
@@ 7.1. Self-Occluded Tap Detection
We first focus on taps and use the common heuristic of X-axis Euler rotation (rotation axis perpendicular to the finger and palm normal) of the metacarpal joint as a measurement of tap completion, which can be used as a continuous value for fine-grained control or a binary “tap” / “no tap” through a simple threshold. We analyze only taps where the finger is over 50% occluded to measure partial self-occlusion performance in HaMeR. We do not evaluate the ring finger due to participant discomfort in the data collection process. When evaluating across our collected tap data, we find that DeltaDorsal significantly reduces the angular RMSE of both the index and middle finger. We attribute this difference to the fact that when tapping the index and middle finger, the entire dorsal region tends to stretch across the knuckle more amplifying our signal. This is further supported as DeltaDorsal performs worse on pinky tap detection (+4.60°) which we attribute to the fact that the pinky induces a far smaller dorsal skin change when tapped. Thus, DeltaDorsal demonstrates better capabilities for fine-grain control in small gestures like taps when the fingers are not totally visible and dorsal deformations are easily isolated like in the index finger, with more limited applications in the pinky which appears robust for coarse event detection but not continuous precision control. We believe that this can generalize to other interaction like surface taps by reducing angular error in microgesture scenarios and even non-occluded cases by using the dorsal region as another prior to predict depth.
@@ 7.2. Self-Occluded Pinch Detection
Following our approach in Section 7.1, we use the common heuristic of finger tip to thumb tip distance as a measurement of pinch completion, which can be used as a continuous value for fine-grained control or binary classification, and analyze at least 50% occluded pinch scenarios. When comparing against HaMeR on our collected pinch data (Table 7), we find that DeltaDorsal significantly outperforms HaMeR in index pinches (Dist. RMSE: 8.65-8.65mm) but but performs worse for both middle and ring finger pinches (Dist. RMSE: +5.105.10mm and +3.093.09mm respectively). We believe this is caused by how people generally pinch, where middle finger and ring finger pinches often induce small dorsal region changes as the thumb moves towards the finger. This phenomenon is reduced in the index finger which we found moves more during pinching.
@@ 8. Discussion
Addressing self-occlusion. Through our motivating study in Section 3, we confirm the impact of self-occlusion on pose estimation performance demonstrated in prior works (Park et al., 2022; Li et al., 2023). Our analysis shows that over 20% of egocentric camera captures are affected by self-occlusion, which can increase pose estimation angular error by over $15$ degrees in SOTA models (Figure 8). This severely limits existing systems, forcing users to keep their fingers in view of the egocentric camera to detect a specific gesture. By utilizing the more visible dorsal area, we are able to mitigate these effects (m=4.59m=-4.59) and enable a more expansive interaction area without the need for new sensors beyond an egocentric camera. Future work should continue to focus on methods to address the impacts of self-occlusion, fusing in temporal pose features and semantic gestural prediction to create a more seamless interaction experience.
@@
Applicability. While our system outperformed SOTA baselines when tested on our lab dataset, more work is needed to understand how our approach can generalize to in-the-wild scenarios. Our LOSO experiments and skin tone study results indicate that leveraging skin deformations can generalize to out-of-distribution subjects and serve as a reliable data source even without model personalization or fine-tuning. Future work should aim to conduct a study with a larger population with a wider range of skin, skin age, hair characteristics, lighting conditions, and blemishes to better understand how the utility of dorsal features can change depending on the population and setting at hand.
@@
Thirdly, we found through our ablation tests (Table 5) that small lightweight backbones like ViT-S+ could reasonably provide usable performance for far less compute and shorter inference time. Future work should continue to investigate this idea, offloading detail capture to high-resolution sensors which can then be processed faster and more efficiently on wearable devices like lightweight AR glasses and more which previously could not support accurate single-camera hand tracking.
Finally, we note that DeltaDorsal requires a high-resolution camera feed to maximize performance and capture dorsal features during motion blur induced by fast movements of the hand and head, requiring power and processing. With the rapid pace of improvement in mobile camera technology however, high quality egocentric head-worn cameras are already starting to be integrated into headsets. For example, Apple Vision Pro (2), Pico 4 (48), and Ray-Ban Meta Glasses (STORE, ) have integrated color cameras from 7 to 16 megapixels to enable applications such as passthrough and photography. We acknowledge that using this hardware for hand tracking will increase power draw and computational complexity to effectively leverage our approach. This is demonstrated in our ablation tests across different image sizes shown in Table 4. Nonetheless, we believe DeltaDorsal can still make a tangible impact in existing systems given that it enables the generalization of the large amounts of monocular egocentric data are currently readily available, especially relevant given recent the recent push for egocentric imitation learning in robotics. DeltaDorsal is also a far smaller model than HaMeR with similar performance across all scenarios, self-occluded or not, allowing for deployment onto more hardware constrained devices that may not be able to to support and run a ViT-H based backbone. Furthermore, we note that our approach can still be applied to low resolution images for low-cost monocular sensing systems as demonstrated in our ablation tests on both image size (Table 4) and backbone size (Table 5). The camera can also be mounted closer to the hand (e.g., on the chest) to improve the resolution of dorsal features without changing camera quality.
@@
Extending beyond hand pose. DeltaDorsal uses changes in dorsal skin induced by the complex tendon activations and tension in the skin to predict pose. These same features can extend beyond hand sensing, enabling grasp and shape detection in held objects as demonstrated in Zhai et al. (Zhai et al., 2023). Beyond just grasp, the same features that were used for isometric click detection can also be used to sense the hardness of held objects, enabling both hand and object detection. Dorsal skin has also historically been used in medical sensing, serving as a strong signal for aging (Gao et al., 2011) and dehydration (Goehring et al., 2022), as well as general tendon usage. By analyzing specifically dorsal activation, we can also build models of biomechanical activation which can enable generic wellness and hand fatigue tracking in long-form egocentric use cases.

### v2

@@ 1. Introduction
Hand tracking is a cornerstone of modern human–computer interaction (Chakravarthi et al., 2023; Ungureanu et al., 2020; Zhang et al., 2020), enabling intuitive gesture-based control in Extended Reality (XR), wearables (Yeo et al., 2019), and other interactive systems (Pei et al., 2022; Mollyn and Harrison, 2024). As these technologies expand, the need for accurate and robust hand pose estimation continues to grow, especially from an egocentric point-of-view. However, a key challenge for egocentric hand pose estimation remains: self-occulusion. From this viewpoint, the palm often partially or fully obscures the very fingers being tracked. Our analysis of prominent egocentric hand datasets (Ohkawa et al., 2023; Grauman et al., 2024b; Fan et al., 2023; Kwon et al., 2021) reveals that a finger is fully occluded in nearly a fifth of all scenarios, with partial occlusions occurring even more frequently.
@@
    A 12-participant evaluation of DeltaDorsal’s performance versus state-of-the-art baselines, as well as analyses with respect to occlusion, skin tone, image size, and backbone.
@@
    Several exemplary use cases of DeltaDorsal in key hand interactions, including detection of pinching, tapping, and isometric force click.
@@ 2.2. Hand Pose Estimation Leveraging Skin Features
Vision-based approaches observing skin features free the dorsum area from on-hand sensors. Opisthenar uses a wrist-worn camera to capture dorsal features for gesture and tap recognition (Yeo et al., 2019). DorsalNet uses the same camera platform with a hand simulator to capture motion images of the dorsum to predict 3D hand poses (Wu et al., 2020). These camera-based methods all require hardware setups on the wrist or arm, making them less usable compared to the more traditional egocentric hand tracking systems.
@@ 3.1. Measuring Self-Occlusion
Using the hand mesh $\Theta$, we project vertices $V$ into the egocentric camera space using the camera extrinsic parameters $V_{\textrm{cam}}=RV_{\textrm{world}}+t$ collected from initial camera calibration. For each face of the MANO model, we compute its normal vector and filter to only faces with a normal vector facing the camera (z-z). We then apply Z-buffering (Catmull, 1978) using this filtered list of faces. The 3D mesh is projected into the camera’s image plane where each mesh face becomes a 2D triangle. We then create an array of depth values, known as a Z-buffer, to track the closest distance to some face at that pixel. For each pixel, we compute the triangle membership through barycentric coordinate sampling and track the interpolated depth at that point. If the interpolated depth is equal or less than the stored Z-buffer depth, the face is marked as visible; otherwise it is occluded and ignored. We then compute the surface area of each visible face as a measurement of mesh visibility. This measurement only considers visibility in respect to other faces of the hand and the camera view. Thus, this only measures the amount of self-occlusion per frame and does not account for occlusion from manipulating other objects or body parts, which disproportionately affects the fingers and palm.
@@ 3.2. Occlusion Prevalence in Egocentric Data
Hand self-occlusion. As shown in Figure 3a, more than 20% of frames across all evaluated datasets exhibit at least one occluded finger, with over 5% containing two or more occluded fingers. These values are a conservative estimate as we do not consider any articulating objects, which are far more likely to occlude the fingers than the dorsum of the hand. These self-occlusion cases often occur because of three primary scenarios: The hand is angled downwards such that the back of the hand occludes the fingers, the hand is angled out such that the index finger occludes the other fingers, or the hand is partially out of frame. Our analysis indicates that there is a sizable proportion of egocentric hand capture data affected by self-occlusion which follows findings from literature. (Fan et al., 2024).
Hand dorsum visibility. We then analyze how much of the dorsum is visible in these situations by computing the distribution of dorsal face visibility with at least one finger over 90% occluded. We find that for every dataset, the median dorsal visibility is above 40% and up to 80%, indicating that a significant proportion of frames with occluded fingers still have visible dorsal features. Thus, dorsal features offer an alternative rich and stable signal when hand and finger silhouettes are not available.
@@ 3.3. Impact of Occlusion on Egocentric Hand Pose Estimators
We find that for both HaMeR and HandOccNet, there is a statistically significant ($p<.0001$) negative correlation between finger visibility and error ($m=-15.84\degree$, $R^{2}=0.212$ for HaMeR, $m=-9.11\degree$, $R^{2}=0.095$ for HandOccNet). This is especially apparent in HaMeR, which has been widely regarded as a standard in pose estimation since its release and has become the backbone of many dataset annotation and hand pose estimation systems (Zhao et al., 2024a; Prakash et al., 2024). Notably for HandOccNet, we found that the model would often default to outputting a neutral hand position in self-occlusion situations, leading to serious prediction errors most likely due to the model’s limited precision in predicting hands occluded behind other objects. Our findings are consistent with our intuition that most existing hand tracking models rely on the general silhouette that lacks the information fidelity to predict pose. Often, their architecture prevents them from using skin features due to the limited token and image size allowed.
@@ 3.4. Our Proposal: Investigate Dorsal Features
Given our findings, we choose to investigate dorsal features as a key signal to improve hand pose estimation in egocentric POVs. We lay out our work as follows: we first perform a high-resolution data collection of dorsal features from different hand gestures to capture and analyze dorsal features in detail. Using this dataset, we propose and train a new model that normalizes visual features from the dorsum to a neutral hand pose, similar to how DorsalNet uses motion images (Wu et al., 2020), to isolate dorsal features even through global motion. We then evaluate our system against baseline hand pose estimation models and across skin tones and image resolutions. Finally, we demonstrate the utility of dorsal features beyond pose estimation by using them in the detection of interaction modalities including pinch, tap, and click.
@@ 5.2. Network Architecture
DeltaDorsal adopts a two-stream architecture built on a transformer backbone with a lightweight convolutional change encoder followed by a regression head. A visualization of this architecture is shown in Figure 6. Using a pretrained DINOv3 Vision Transformer (ViT) (Siméoni et al., 2025) following the “large” design (ViT-L) with the last three blocks unfrozen, we first compute the dense image features for both the reference $I_{0}$ and target $I_{t}$ using the image of the dorsum of the hand, $F_{0}$ and $F_{t}$ respectively with dimensions equal to the original image size divided by the patch size ($16\times 16$). Notably, model weights are shared across each stream to ensure that $F_{0}$ and $F_{t}$ maintain the same feature space. A change encoder then processes this feature map by fusing the feature delta, $F_{t}-F_{0}$, cosine similarity per patch, $\cos(F_{t},F_{0})(x,y)$, $F_{t}$, and $F_{0}$ and feeding this data through a small convolutional network to generate the encoded features $X$. Cosine similarity features are visualized in Figure 5. By fusing the original features, feature deltas, and cosine similarities in the change encoder, we allow the model to select between general image features and changes in dorsal features as needed. We settled on a dense featurizer backbone after preliminary experiments with optical flow methods like RAFT (Teed and Deng, 2020) and Gunnar-Farneback optical flow (Farnebäck, 2003) had difficulty separating skin deformations from global movements of the hand. Furthermore, traditional feature extractors like ResNet (He et al., 2015) and ConvNeXt (Liu et al., 2022) failed to generalize to out-of-distribution subjects. We believe this limitation arises because these models are not well-suited to capture the subtle soft-body surface deformations in the back of the hand which often do not demonstrate distinct, localized feature changes. Finally, a regression head applies a small convolutional neck for local refinement, pools spatial features, and passes pooled vectors through a shallow multi-layer perceptron to predict the pose vector $\beta$ ($(15,3)$ for 15 axis-angle rotations).
@@ 5.3. Implementation Details
All systems were implemented using PyTorch (Paszke et al., 2019). We used the AdamW optimizer (Loshchilov and Hutter, 2019) with a batch size of 96 for our training. During training, we randomly selected a reference frame from the corresponding calibration data capture. Please refer to our codebase and configuration files for specific parameter settings for experiments. All training and evaluations were performed using PyTorch Distributed Data Parallel on a system with an AMD Ryzen Threadripper PRO 3955WX 3.90GHz, four Nvidia RTX A5500s, and 256GB of RAM. Our full codebase can be found at https://github.com/hilab-open-source/deltadorsal.
@@ 6.1. Pose Estimation Performance Study
Performance across occlusion levels. We also evaluate the performance of our baselines and system under different occlusion levels after fine-tuning. As shown in Figure 8, we find that our model has a smaller dropoff in performance compared to both HaMeR and HandOccNet ($m=-4.59\degree$, $R^{2}=0.048$ for ours, $m=-8.42\degree$, $R^{2}=0.117$ for HaMeR, $m=-9.11\degree$, $R^{2}=0.095$ for HandOccNet) [^1] We denote slope as $m$. HandOccNet’s general performance was not significantly changed after finetuning due to previously mentioned training difficulties. When comparing performance at 50% finger visibility and lower, DeltaDorsal outperforms HaMeR across both MPJAE (Ours: $7.59\pm 4.38\degree$, HaMeR: $9.24\pm 5.54\degree$, percent change: $17.86\%$) and PA-MPJPE (Ours: $8.47\pm 3.84$mm, HaMeR: $9.03\pm 3.65$mm, percent change: $6.20\%$). These findings are consistent with our intuition that existing hand pose estimation models rely heavily on finger silhouettes for pose prediction. Examples of this poor performance are shown in Figure 1 and Figure 7, where HaMeR and HandOccNet both fail to accurately predict hand pose when the fingers are occluded.
@@ 7.1. Self-Occluded Tap Detection
We first focus on taps and use the common heuristic of X-axis Euler rotation (rotation axis perpendicular to the finger and palm normal) of the metacarpal joint as a measurement of tap completion, which can be used as a continuous value for fine-grained control or a binary “tap” / “no tap” through a simple threshold. We analyze only taps where the finger is over 50% occluded to measure partial self-occlusion performance in HaMeR. We do not evaluate the ring finger due to participant discomfort in the data collection process. When evaluating across our collected tap data, we find that DeltaDorsal significantly reduces the angular RMSE of both the index and middle finger. We attribute this difference to the fact that when tapping the index and middle finger, the entire dorsal region tends to stretch across the knuckle more amplifying our signal. This is further supported as DeltaDorsal performs worse on pinky tap detection (+4.60°) which we attribute to the fact that the pinky induces a far smaller dorsal skin change when tapped. Thus, DeltaDorsal demonstrates better capabilities for fine-grain control in small gestures like taps when the fingers are not totally visible and dorsal deformations are easily isolated like in the index finger, with more limited applications in the pinky which appears robust for coarse event detection but not continuous precision control. We believe that this can generalize to other interactions like surface taps by reducing angular error in microgesture scenarios and even non-occluded cases by using the dorsal region as another prior to predict depth.
@@ 7.2. Self-Occluded Pinch Detection
Following our approach in Section 7.1, we use the common heuristic of finger tip to thumb tip distance as a measurement of pinch completion, which can be used as a continuous value for fine-grained control or binary classification, and analyze at least 50% occluded pinch scenarios. When comparing against HaMeR on our collected pinch data (Table 7), we find that DeltaDorsal significantly outperforms HaMeR in index pinches (Dist. RMSE: 8.65-8.65mm) but performs worse for both middle and ring finger pinches (Dist. RMSE: +5.105.10mm and +3.093.09mm respectively). We believe this is caused by how people generally pinch, where middle finger and ring finger pinches often induce small dorsal region changes as the thumb moves towards the finger. This phenomenon is reduced in the index finger which we found moves more during pinching.
@@ 8. Discussion
Addressing self-occlusion. Through our motivating study in Section 3, we confirm the impact of self-occlusion on pose estimation performance demonstrated in prior works (Park et al., 2022; Li et al., 2023). Our analysis shows that over 20% of egocentric camera captures are affected by self-occlusion, which can increase pose estimation angular error by over $15$ degrees in SOTA models (Figure 8). This severely limits existing systems, forcing users to keep their fingers in view of the egocentric camera to detect a specific gesture. By utilizing the more visible dorsal area, we are able to mitigate these effects and enable a more expansive interaction area without the need for new sensors beyond an egocentric camera. Future work should continue to focus on methods to address the impacts of self-occlusion, fusing in temporal pose features and semantic gestural prediction to create a more seamless interaction experience.
@@
Generalizability. While our system outperformed SOTA baselines when tested on our lab dataset, more work is needed to understand how our approach can generalize to in-the-wild scenarios. Our LOSO experiments and skin tone study results indicate that leveraging skin deformations can generalize to out-of-distribution subjects and serve as a reliable data source even without model personalization or fine-tuning. Future work should aim to conduct a study with a larger population with a wider range of skin, skin age, hair characteristics, lighting conditions, and blemishes to better understand how the utility of dorsal features can change depending on the population and setting at hand.
@@
Deployability. We found through our ablation tests (Table 5) that small lightweight backbones like ViT-S+ could reasonably provide usable performance for far less compute and shorter inference time. Future work should continue to investigate this idea, offloading detail capture to high-resolution sensors which can then be processed faster and more efficiently on wearable devices like lightweight AR glasses and more which previously could not support accurate single-camera hand tracking.
We also note that DeltaDorsal requires a high-resolution camera feed to maximize performance and capture dorsal features during motion blur induced by fast movements of the hand and head, requiring power and processing. However, high quality egocentric head-worn cameras are already starting to be integrated into headsets (2; 48; 58). We acknowledge that using this hardware for hand tracking will likely increase power draw and computational complexity. Nonetheless, we believe DeltaDorsal can still make a tangible impact in existing systems given that it enables the utilization of the large amounts of monocular egocentric data are currently readily available.
DeltaDorsal is also a far smaller model than HaMeR with similar performance across all scenarios, self-occluded or not, allowing for deployment onto more hardware constrained devices that may not be able to to support and run a ViT-H based backbone.
Finally, we note that our approach can still be applied to low resolution images for low-cost monocular sensing systems as demonstrated in our ablation tests on both image size (Table 4) and backbone size (Table 5). The camera can also be mounted closer to the hand (e.g., on the chest) to improve the resolution of dorsal features without changing camera quality.
@@
Beyond hand pose estimation. DeltaDorsal uses changes in dorsal skin induced by the complex tendon activations and tension in the skin to predict pose. These same features can extend beyond hand sensing, enabling grasp and shape detection in held objects as demonstrated in Zhai et al. (Zhai et al., 2023). Beyond just grasp, the same features that were used for isometric click detection can also be used to sense the hardness of held objects, enabling both hand and object detection. Dorsal skin has also historically been used in medical sensing, serving as a strong signal for aging (Gao et al., 2011) and dehydration (Goehring et al., 2022), as well as general tendon usage. By analyzing specifically dorsal activation, we can also build models of biomechanical activation which can enable generic wellness and hand fatigue tracking in long-form egocentric use cases.